# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 10.1.2 Parametric Demand
Broder and Rusmevichientong (2012) consider a parametric demand model $d(p;z)=\Pr(V\geq p)$, where $z\in\mathbb{R}^{n}$ is an unknown parameter governing the demand curve. The seller observes binary purchase decisions and updates a maximum likelihood estimate of $z$. Under standard regularity conditions (bounded demand, unique optimal price, smooth revenue function), the minimax regret is again $\Theta(\sqrt{T})$.179179179The lower bound (Theorem 3.1 of Broder and Rusmevichientong (2012)) constructs a linear demand family where all demand curves pass through the same point at the optimal price $p^{*}(z_{0})$. Observing purchases at this price provides no information about $z$. An MLE-Cycle policy that interleaves dedicated exploration rounds with greedy pricing achieves the matching upper bound $O(\sqrt{T})$ (Theorem 3.6). Parametric structure alone does not break the $\sqrt{T}$ barrier.
The picture changes under a “well-separated” condition requiring that every price in the feasible set is informative about the demand parameter. Formally, the Fisher information $I(p,z)$ is bounded below by $c_{f}>0$ for all prices $p$ and parameters zz.180180180This means no two demand curves $d(p;z)$ and $d(p;z^{\prime})$ cross on the pricing interval, so every purchase observation distinguishes the two hypotheses. In the lower-bound family of Theorem 3.1, the curves all cross at $p=1$, which is why the $\sqrt{T}$ floor emerges. Under well-separation, a pure greedy policy (MLE-Greedy, pricing at $p^{*}(\hat{z}_{t})$ each period) achieves $O(\log T)$ regret (Theorem 4.8), because every price is informative and dedicated exploration rounds become unnecessary.
@@ 10.1.3 High-Dimensional Features with Sparsity
Javanmard and Nazerzadeh (2019) extend the setting to products described by $d$-dimensional feature vectors $x_{t}$. The market value is $v_{t}=x_{t}^{\top}\theta_{0}+z_{t}$, where $\theta_{0}\in\mathbb{R}^{d}$ is unknown and $z_{t}$ is i.i.d. noise from a known log-concave distribution FF.181181181Log-concavity of $F$ and 1F1-F is satisfied by the normal, logistic, uniform, Laplace, and exponential distributions. It ensures that expected revenue $p\cdot[1-F(p-x_{t}^{\top}\theta_{0})]$ is strictly quasi-concave in $p$, giving a unique optimal price. Only $s_{0}$ of the $d$ coordinates of $\theta_{0}$ are nonzero, but the seller does not know which ones.
Their RMLP algorithm (Regularized Maximum Likelihood Pricing) operates in episodes whose lengths double (1, 2, 4, 8, …). At each episode boundary, the seller fits a LASSO-penalized maximum likelihood estimate of $\theta_{0}$ using data from the previous episode, then prices greedily throughout the current episode. The regret is $O(s_{0}\log d\cdot\log T)$ (Theorem 4.1), with a matching lower bound of $\Omega(s_{0}(\log d+\log T))$ (Theorem 5.1). The reason this beats $\sqrt{T}$ connects back to the Broder lower bound. In the parametric model of Section 10.1.2, all demand curves can cross at the optimal price, making that price uninformative. Here, customer features vary across periods, so the aggregate demand function at any fixed price changes in proportion to the estimation error in $\theta_{0}$. Every price is informative, and dedicated exploration rounds become unnecessary.182182182If some feature directions are rarely observed, the seller cannot learn all coordinates of $\theta_{0}$ quickly, and regret degrades to $O(\sqrt{\log(d)\cdot T})$ (Theorem 4.2). If the noise distribution belongs to a known parametric family but its scale parameter is unknown, regret reverts to $\Omega(\sqrt{T})$ (Theorem 7.1), foreshadowing the result of Xu and Wang (2021) discussed in Section 10.3. Even with hundreds of features, if only a handful matter, learning is fast.
@@ 10.2 Revealed Preference and Partial Identification
if $p_{k}$ is not dominated, and $I_{p_{k}}(t)=0$ otherwise, where $\hat{\mu}_{p_{k}}(t)$ is the average profit observed at price $p_{k}$ and $N_{p_{k}}(t)$ is the number of trials.183183183Standard UCB1 uses an exploration bonus of $\sqrt{2\ln t/N_{p_{k}}(t)}$, which assumes rewards in $[0,1]$. Since profit at price $p_{k}$ is bounded by $p_{k}$, scaling the bonus by $p_{k}$ tightens exploration for cheap prices that cannot contribute much profit regardless. Two innovations drive the improvement over standard UCB1. First, dominance elimination reduces the effective number of arms the algorithm must explore. Second, the $p_{k}$ scaling focuses exploration on prices where uncertainty actually matters for profit. Together, these yield $O(\log T)$ regret.
@@
where $\Delta_{k}=\mu_{k^{*}}-\mu_{k}$ is the gap between arm $k$ and the optimal arm.184184184The first sum is the leading term, scaling as $O(\log T)$. The second sum is a constant that does not grow with $T$. Replacing $p_{k}$ with 1 recovers the standard UCB1 bound, which is looser.
Misra et al. (2019) calibrate the model to a field experiment at ZipRecruiter, a B2B online recruiting platform. With 7,870 customers per month, 1,000 segments, and 10 price points from \$19 to \$399, UCB-PI achieves 98% of oracle profit and produces 43% higher profits during the first month of testing compared to a learn-then-earn alternative. The algorithm has both higher mean profit and lower variance, reflecting the value of eliminating dominated prices early.185185185For multi-product settings, Mueller et al. (2019) impose low-rank structure on the price-sensitivity matrix, achieving regret $O(T^{3/4}\sqrt{d})$ that scales with the latent demand dimension $d$ rather than the number of products. Badanidiyuru et al. (2013) extend the bandit framework to handle inventory constraints (“bandits with knapsacks”), relevant when the seller faces limited stock alongside the pricing decision.
@@ 10.3 The Value of Knowing the Noise Distribution
Xu and Wang (2021) ask how much it helps to know the shape of demand uncertainty. In their model, a feature vector $x_{t}\in\mathbb{R}^{d}$ describes each sales session, the customer’s valuation is $w_{t}=x_{t}^{\top}\theta^{*}+N_{t}$ where $\theta^{*}$ is unknown and $N_{t}$ is zero-mean i.i.d. noise with CDF $F$, and the seller observes only whether the customer bought at the posted price. The answer depends on whether $F$ is known.186186186The regret benchmark here differs from Section 10.1.1. Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) measure regret against the best fixed price; Xu and Wang (2021) and Liu et al. (2024) measure regret against the clairvoyant contextual policy that sets the optimal price $p_{t}^{*}$ for each customer’s features $x_{t}$. The contextual benchmark is harder, which makes the $O(d\log T)$ rate more impressive.
@@
If $F$ is known (the seller knows that demand shocks are, say, normally distributed with known variance), the EMLP algorithm (Epoch-based Maximum Likelihood Pricing) achieves regret $O(d\log T)$ (Theorem 3).187187187EMLP runs in doubling epochs of length $\tau_{k}=2^{k-1}$. At each epoch boundary, the seller fits a maximum likelihood estimate $\hat{\theta}_{k}$ using data from the previous epoch, then prices greedily at $p_{t}=J(x_{t}^{\top}\hat{\theta}_{k})$ throughout the epoch, where $J(u)=\arg\max_{v}\,v[1-F(v-u)]$ is the revenue-maximizing price function. The key technical insight is that the negative log-likelihood is strongly convex (Lemma 7 of Xu and Wang (2021)), so MLE concentrates at rate $O(d/\tau_{k})$. Since regret is quadratic in the parameter estimation error (Lemma 5) and there are $O(\log T)$ epochs, the total regret is $O(d\log T)$. After 10,000 customers with $d=5$ features, the revenue loss is roughly 50 customers’ worth, an improvement over the $\sqrt{T}\approx 100$ customers’ worth that Kleinberg’s lower bound imposes without structural knowledge.
@@
If $F$ is unknown (even if only the variance of a Gaussian is unknown, with everything else known), the regret is at least $\Omega(\sqrt{T})$ (Theorem 12).188188188The lower bound constructs two noise variances $\sigma_{1}=1$ and $\sigma_{2}=1-T^{-1/4}$. Any algorithm that performs well under both must spend $\Omega(\sqrt{T})$ revenue distinguishing the two cases. This extends the “uninformative price” phenomenon of Broder and Rusmevichientong (2012): when the seller does not know $F$, there exist prices at which observed purchase behavior is nearly identical under different demand parameters. The seller is back to the Kleinberg baseline, with no algorithm able to achieve sublinear improvement regardless of how many features are available.
@@
The gap between $O(d\log T)$ and $\Omega(\sqrt{T})$ is super-polynomial in $T$, not merely a constant factor. When an economist specifies a logit or probit demand model, the assumed noise distribution purchases a qualitative improvement in learning rate. Semiparametric approaches that leave the error distribution unspecified pay a concrete cost, reverting from logarithmic to polynomial regret.189189189Tullii et al. (2024) establish the tightest known bound under minimal distributional assumptions: if the noise distribution (c.d.f.) is merely Lipschitz continuous, the minimax regret is $\Theta(T^{2/3})$, strictly between the $\log T$ rate with known $F$ and the $\sqrt{T}$ rate with unknown $F$. Fan et al. (2024) consider a semiparametric setting where the noise density is smooth and connect the pricing problem to the econometrics of semiparametric estimation.
@@ 10.4 Strategic Buyers
where $A$ is a positive definite matrix governing manipulation costs.190190190The matrix $A$ captures how costly it is for the buyer to distort each feature dimension. High eigenvalues of $A$ mean manipulation is expensive. This is the standard model of strategic classification (Hardt et al., 2016), adapted to pricing. The first-order condition yields a systematic bias: the buyer shifts her features to make the seller’s model predict a lower valuation, securing a lower price. The seller observes only the distorted features $\tilde{x}_{t}$, not the true $x_{t}^{0}$.
@@
Incentive compatibility matters even in settings where the seller is “just” learning demand.191191191Agrawal and Tang (2024) document a related phenomenon in pricing with reference effects. If consumers anchor on past prices, a static pricing policy that ignores reference dependence accumulates linear regret $\Omega(T)$. Chen et al. (2025) show that imposing fairness constraints (requiring similar prices for similar customers) raises the regret floor to $\Theta(T^{2/3})$, a social cost of equitable treatment.
@@ 10.7 Simulation Study: The Knowledge Ladder
I run six algorithms on the Misra et al. (2019) demand environment to trace how cumulative regret responds to increasing structural knowledge.192192192The environment has $K=100$ prices on a grid from \$0.01 to \$1.00, S=1, [^0] S=1{,}000 consumer segments with equal weights, within-segment heterogeneity $\delta=0.1$, and segment midpoints $v_{s}\sim\mathrm{Uniform}(0.1,0.9)$. A consumer purchases if and only if $v_{i}\geq p$ (WARP). Each algorithm runs across 10 seeds with T=200, [^0] T=200{,}000 rounds. The six algorithms, ordered by the structural knowledge they exploit, are: (0) $\varepsilon$-greedy with $\varepsilon=0.1$, which never adapts its exploration rate; (1) Learn-Then-Earn (LTE), which explores uniformly for the first 5% of rounds and then commits to the empirical best price; (2) UCB1 (Auer et al., 2002a), which adapts exploration via confidence bounds but ignores demand structure; (3) Thompson Sampling (Thompson, 1933), which maintains a Bayesian posterior over purchase rates193193193At each period the algorithm draws one sample $\tilde{\mu}_{k}$ from each arm’s posterior over its purchase rate and selects the arm with the highest sampled expected profit $p_{k}\tilde{\mu}_{k}$; arms with uncertain posteriors have high-variance draws and are selected frequently, while well-estimated arms are selected in proportion to how likely they are optimal.; (4) UCB-PI (Misra et al., 2019), which uses WARP to eliminate dominated prices and scales the exploration bonus by the price level; and (5) UCB-PI-tuned, which adds a variance-based refinement to the exploration bonus.
@@ 11 Offline Reinforcement Learning and Human Feedback
This is the problem of offline reinforcement learning, also called batch reinforcement learning.194194194The term “batch RL” was standard in the earlier literature (Ernst et al., 2005; Lange et al., 2012). “Offline RL” became dominant after Levine et al. (2020), who distinguished it from off-policy RL (which still collects new data, just under a different policy than the target). I use “offline RL” throughout. The agent never queries the environment. All learning happens from a fixed dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i},s_{i}^{\prime})\}_{i=1}^{n}$ collected by some behavioral policy $\pi_{b}$. The goal is to find a policy $\hat{\pi}$ whose value $V^{\hat{\pi}}$ is as close to the optimal $V^{*}$ as possible.
Online RL algorithms (Q-learning, SARSA, policy gradient methods from Section 4) can in principle be applied to offline data by treating the dataset as a replay buffer. In practice, this fails catastrophically. The reason is distributional shift: the learned policy $\hat{\pi}$ inevitably queries state-action pairs that the behavioral policy $\pi_{b}$ never visited, and the Q-function at these unseen pairs is pure extrapolation. Because the Bellman backup propagates these extrapolation errors through the max operator, errors compound geometrically across the planning horizon, producing arbitrarily poor policies.195195195This failure mode was first demonstrated empirically by Fujimoto et al. (2019), who showed that standard off-policy algorithms (DDPG, SAC) trained purely from a static dataset performed worse than the behavioral policy itself, even when that behavioral policy was a partially-trained, mediocre agent. The gap widened with dataset size, the opposite of what one expects from more data.

### 23 Mar 2026 v4

@@ 10.1.2 Parametric Demand
Broder and Rusmevichientong (2012) consider a parametric demand model $d(p;z)=\Pr(V\geq p)$, where $z\in\mathbb{R}^{n}$ is an unknown parameter governing the demand curve. The seller observes binary purchase decisions and updates a maximum likelihood estimate of $z$. Under standard regularity conditions (bounded demand, unique optimal price, smooth revenue function), the minimax regret is again $\Theta(\sqrt{T})$.189189189The lower bound (Theorem 3.1 of Broder and Rusmevichientong (2012)) constructs a linear demand family where all demand curves pass through the same point at the optimal price $p^{*}(z_{0})$. Observing purchases at this price provides no information about $z$. An MLE-Cycle policy that interleaves dedicated exploration rounds with greedy pricing achieves the matching upper bound $O(\sqrt{T})$ (Theorem 3.6). Parametric structure alone does not break the $\sqrt{T}$ barrier.
The picture changes under a “well-separated” condition requiring that every price in the feasible set is informative about the demand parameter. Formally, the Fisher information $I(p,z)$ is bounded below by $c_{f}>0$ for all prices $p$ and parameters zz.190190190This means no two demand curves $d(p;z)$ and $d(p;z^{\prime})$ cross on the pricing interval, so every purchase observation distinguishes the two hypotheses. In the lower-bound family of Theorem 3.1, the curves all cross at $p=1$, which is why the $\sqrt{T}$ floor emerges. Under well-separation, a pure greedy policy (MLE-Greedy, pricing at $p^{*}(\hat{z}_{t})$ each period) achieves $O(\log T)$ regret (Theorem 4.8), because every price is informative and dedicated exploration rounds become unnecessary.
@@ 10.1.3 High-Dimensional Features with Sparsity
Javanmard and Nazerzadeh (2019) extend the setting to products described by $d$-dimensional feature vectors $x_{t}$. The market value is $v_{t}=x_{t}^{\top}\theta_{0}+z_{t}$, where $\theta_{0}\in\mathbb{R}^{d}$ is unknown and $z_{t}$ is i.i.d. noise from a known log-concave distribution FF.191191191Log-concavity of $F$ and 1F1-F is satisfied by the normal, logistic, uniform, Laplace, and exponential distributions. It ensures that expected revenue $p\cdot[1-F(p-x_{t}^{\top}\theta_{0})]$ is strictly quasi-concave in $p$, giving a unique optimal price. Only $s_{0}$ of the $d$ coordinates of $\theta_{0}$ are nonzero, but the seller does not know which ones.
Their RMLP algorithm (Regularized Maximum Likelihood Pricing) operates in episodes whose lengths double (1, 2, 4, 8, …). At each episode boundary, the seller fits a LASSO-penalized maximum likelihood estimate of $\theta_{0}$ using data from the previous episode, then prices greedily throughout the current episode. The regret is $O(s_{0}\log d\cdot\log T)$ (Theorem 4.1), with a matching lower bound of $\Omega(s_{0}(\log d+\log T))$ (Theorem 5.1). The reason this beats $\sqrt{T}$ connects back to the Broder lower bound. In the parametric model of Section 10.1.2, all demand curves can cross at the optimal price, making that price uninformative. Here, customer features vary across periods, so the aggregate demand function at any fixed price changes in proportion to the estimation error in $\theta_{0}$. Every price is informative, and dedicated exploration rounds become unnecessary.192192192If some feature directions are rarely observed, the seller cannot learn all coordinates of $\theta_{0}$ quickly, and regret degrades to $O(\sqrt{\log(d)\cdot T})$ (Theorem 4.2). If the noise distribution belongs to a known parametric family but its scale parameter is unknown, regret reverts to $\Omega(\sqrt{T})$ (Theorem 7.1), foreshadowing the result of Xu and Wang (2021) discussed in Section 10.3. Even with hundreds of features, if only a handful matter, learning is fast.
@@ 10.2 Revealed Preference and Partial Identification
if $p_{k}$ is not dominated, and $I_{p_{k}}(t)=0$ otherwise, where $\hat{\mu}_{p_{k}}(t)$ is the average profit observed at price $p_{k}$ and $N_{p_{k}}(t)$ is the number of trials.193193193Standard UCB1 uses an exploration bonus of $\sqrt{2\ln t/N_{p_{k}}(t)}$, which assumes rewards in $[0,1]$. Since profit at price $p_{k}$ is bounded by $p_{k}$, scaling the bonus by $p_{k}$ tightens exploration for cheap prices that cannot contribute much profit regardless. Two innovations drive the improvement over standard UCB1. First, dominance elimination reduces the effective number of arms the algorithm must explore. Second, the $p_{k}$ scaling focuses exploration on prices where uncertainty actually matters for profit. Together, these yield $O(\log T)$ regret.
@@
where $\Delta_{k}=\mu_{k^{*}}-\mu_{k}$ is the gap between arm $k$ and the optimal arm.194194194The first sum is the leading term, scaling as $O(\log T)$. The second sum is a constant that does not grow with $T$. Replacing $p_{k}$ with 1 recovers the standard UCB1 bound, which is looser.
Misra et al. (2019) calibrate the model to a field experiment at ZipRecruiter, a B2B online recruiting platform. With 7,870 customers per month, 1,000 segments, and 10 price points from \$19 to \$399, UCB-PI achieves 98% of oracle profit and produces 43% higher profits during the first month of testing compared to a learn-then-earn alternative. The algorithm has both higher mean profit and lower variance, reflecting the value of eliminating dominated prices early.195195195For multi-product settings, Mueller et al. (2019) impose low-rank structure on the price-sensitivity matrix, achieving regret $O(T^{3/4}\sqrt{d})$ that scales with the latent demand dimension $d$ rather than the number of products. Badanidiyuru et al. (2013) extend the bandit framework to handle inventory constraints (“bandits with knapsacks”), relevant when the seller faces limited stock alongside the pricing decision.
@@ 10.3 The Value of Knowing the Noise Distribution
Xu and Wang (2021) ask how much it helps to know the shape of demand uncertainty. In their model, a feature vector $x_{t}\in\mathbb{R}^{d}$ describes each sales session, the customer’s valuation is $w_{t}=x_{t}^{\top}\theta^{*}+N_{t}$ where $\theta^{*}$ is unknown and $N_{t}$ is zero-mean i.i.d. noise with CDF $F$, and the seller observes only whether the customer bought at the posted price. The answer depends on whether $F$ is known.196196196The regret benchmark here differs from Section 10.1.1. Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) measure regret against the best fixed price; Xu and Wang (2021) and Liu et al. (2024) measure regret against the clairvoyant contextual policy that sets the optimal price $p_{t}^{*}$ for each customer’s features $x_{t}$. The contextual benchmark is harder, which makes the $O(d\log T)$ rate more impressive.
@@
If $F$ is known (the seller knows that demand shocks are, say, normally distributed with known variance), the EMLP algorithm (Epoch-based Maximum Likelihood Pricing) achieves regret $O(d\log T)$ (Theorem 3).197197197EMLP runs in doubling epochs of length $\tau_{k}=2^{k-1}$. At each epoch boundary, the seller fits a maximum likelihood estimate $\hat{\theta}_{k}$ using data from the previous epoch, then prices greedily at $p_{t}=J(x_{t}^{\top}\hat{\theta}_{k})$ throughout the epoch, where $J(u)=\arg\max_{v}\,v[1-F(v-u)]$ is the revenue-maximizing price function. The key technical insight is that the negative log-likelihood is strongly convex (Lemma 7 of Xu and Wang (2021)), so MLE concentrates at rate $O(d/\tau_{k})$. Since regret is quadratic in the parameter estimation error (Lemma 5) and there are $O(\log T)$ epochs, the total regret is $O(d\log T)$. After 10,000 customers with $d=5$ features, the revenue loss is roughly 50 customers’ worth, an improvement over the $\sqrt{T}\approx 100$ customers’ worth that Kleinberg’s lower bound imposes without structural knowledge.
@@
If $F$ is unknown (even if only the variance of a Gaussian is unknown, with everything else known), the regret is at least $\Omega(\sqrt{T})$ (Theorem 12).198198198The lower bound constructs two noise variances $\sigma_{1}=1$ and $\sigma_{2}=1-T^{-1/4}$. Any algorithm that performs well under both must spend $\Omega(\sqrt{T})$ revenue distinguishing the two cases. This extends the “uninformative price” phenomenon of Broder and Rusmevichientong (2012): when the seller does not know $F$, there exist prices at which observed purchase behavior is nearly identical under different demand parameters. The seller is back to the Kleinberg baseline, with no algorithm able to achieve sublinear improvement regardless of how many features are available.
@@
The gap between $O(d\log T)$ and $\Omega(\sqrt{T})$ is super-polynomial in $T$, not merely a constant factor. When a modeler specifies a logit or probit demand model, the assumed noise distribution purchases a qualitative improvement in learning rate. Semiparametric approaches that leave the error distribution unspecified pay a concrete cost, reverting from logarithmic to polynomial regret.199199199Tullii et al. (2024) establish the tightest known bound under minimal distributional assumptions: if the noise distribution (c.d.f.) is merely Lipschitz continuous, the minimax regret is $\Theta(T^{2/3})$, strictly between the $\log T$ rate with known $F$ and the $\sqrt{T}$ rate with unknown $F$. Fan et al. (2024) consider a semiparametric setting where the noise density is smooth and connect the pricing problem to the econometrics of semiparametric estimation.
@@ 10.4 Strategic Buyers
where $A$ is a positive definite matrix governing manipulation costs.200200200The matrix $A$ captures how costly it is for the buyer to distort each feature dimension. High eigenvalues of $A$ mean manipulation is expensive. This is the standard model of strategic classification (Hardt et al., 2016), adapted to pricing. The first-order condition yields a systematic bias: the buyer shifts her features to make the seller’s model predict a lower valuation, securing a lower price. The seller observes only the distorted features $\tilde{x}_{t}$, not the true $x_{t}^{0}$.
@@
Incentive compatibility matters even in settings where the seller is “just” learning demand.201201201Agrawal and Tang (2024) document a related phenomenon in pricing with reference effects. If consumers anchor on past prices, a static pricing policy that ignores reference dependence accumulates linear regret $\Omega(T)$. Chen et al. (2025) show that imposing fairness constraints (requiring similar prices for similar customers) raises the regret floor to $\Theta(T^{2/3})$, a social cost of equitable treatment.
@@ 10.7 Simulation Study: The Knowledge Ladder
I run six algorithms on the Misra et al. (2019) demand environment to trace how cumulative regret responds to increasing structural knowledge.202202202The environment has $K=100$ prices on a grid from \$0.01 to \$1.00, S=1, [^0] S=1{,}000 consumer segments with equal weights, within-segment heterogeneity $\delta=0.1$, and segment midpoints $v_{s}\sim\mathrm{Uniform}(0.1,0.9)$. A consumer purchases if and only if $v_{i}\geq p$ (WARP). Each algorithm runs across 10 seeds with T=200, [^0] T=200{,}000 rounds. The six algorithms, ordered by the structural knowledge they exploit, are: (0) $\varepsilon$-greedy with $\varepsilon=0.1$, which never adapts its exploration rate; (1) Learn-Then-Earn (LTE), which explores uniformly for the first 5% of rounds and then commits to the empirical best price; (2) UCB1 (Auer et al., 2002a), which adapts exploration via confidence bounds but ignores demand structure; (3) Thompson Sampling (Thompson, 1933), which maintains a Bayesian posterior over purchase rates203203203At each period the algorithm draws one sample $\tilde{\mu}_{k}$ from each arm’s posterior over its purchase rate and selects the arm with the highest sampled expected profit $p_{k}\tilde{\mu}_{k}$; arms with uncertain posteriors have high-variance draws and are selected frequently, while well-estimated arms are selected in proportion to how likely they are optimal.; (4) UCB-PI (Misra et al., 2019), which uses WARP to eliminate dominated prices and scales the exploration bonus by the price level; and (5) UCB-PI-tuned, which adds a variance-based refinement to the exploration bonus.
@@ 11 Offline Reinforcement Learning and Human Feedback
This is the problem of offline reinforcement learning, also called batch reinforcement learning.204204204The term “batch RL” was standard in the earlier literature (Ernst et al., 2005; Lange et al., 2012). “Offline RL” became dominant after Levine et al. (2020), who distinguished it from off-policy RL (which still collects new data, just under a different policy than the target). I use “offline RL” throughout. The agent never queries the environment. All learning happens from a fixed dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i},s_{i}^{\prime})\}_{i=1}^{n}$ collected by some behavioral policy $\pi_{b}$. The goal is to find a policy $\hat{\pi}$ whose value $V^{\hat{\pi}}$ is as close to the optimal $V^{*}$ as possible.
Online RL algorithms (Q-learning, SARSA, policy gradient methods from Section 4) can in principle be applied to offline data by treating the dataset as a replay buffer. In practice, this fails catastrophically. The reason is distributional shift: the learned policy $\hat{\pi}$ inevitably queries state-action pairs that the behavioral policy $\pi_{b}$ never visited, and the Q-function at these unseen pairs is pure extrapolation. Because the Bellman backup propagates these extrapolation errors through the max operator, errors compound geometrically across the planning horizon, producing arbitrarily poor policies.205205205This failure mode was first demonstrated empirically by Fujimoto et al. (2019), who showed that standard off-policy algorithms (DDPG, SAC) trained purely from a static dataset performed worse than the behavioral policy itself, even when that behavioral policy was a partially-trained, mediocre agent. The gap widened with dataset size, the opposite of what one expects from more data.